# Throwing Vines at the Wall: Structure Learning via Random Search

### v1

@@ 6 CONCLUSION
We propose a simple and yet effective method for learning vine copula structures based on random search and model confidence sets. Our experiments on real-world data sets demonstrate that the new method outperforms existing greedy heuristics in both density estimation and regression tasks, often by a large margin. A limitation of the new method is the additional computational cost compared to greedy heuristics. However, this cost is modest if the dimension is not too high and easily parallelizable, making it a worthwhile trade-off for the significant performance gains achieved. Promising future research directions include improving the search strategy, for example by using more sophisticated sampling methods or prior knowledge about the data, as well as incorporating sparsity constraints for very high-dimensional problems.
@@ Appendix A DETAILS ON THE MCS ALGORITHM
In the following, we recall the relevant details of the MCS algorithms given by Kim and Ramdas, (2025) to keep our paper self-contained. Recall the notation of our main paper, specifically Algorithm 2: There are $M$ candidate models $\{\mathcal{V}_{m}\}_{m=1}^{M}$ and we have already computed validation losses $L_{i,m}$ for $1\leq i\leq n_{\text{val}}$ and $1\leq m\leq M$. For ease of notation, define the loss differences $\Delta_{m,k}^{(i)}=L_{i,m}-L_{i,k}$ for $1\leq i\leq n_{\text{val}}$ and $1\leq m,k\leq M$.
A.1 The DA Test Statistic
We first describe the discrete argmin (DA) hypothesis test introduced by Kim and Ramdas, (2025). This test serves as the foundation for constructing MCS with either pointwise or uniform coverage. The basic algorithm to compute the test statistic from a set of loss observations is given in Algorithm A.3.
@@ A.1 The DA Test Statistic
A.2 Inverting the DA Test for Marginal MCS Construction: the DA-MCS-marg procedure
To obtain an MCS with marginal coverage guarantees, one can simply invert the DA test: compute $\widehat{T}_{r}$ for all $r=1,\dots,M$ by running Algorithm A.3 and define
$\widehat{\Theta}_{marg}=\left\{r\in\{1,\dots,M\}\colon\;\widehat{T}_{r}\leq\Phi^{-1}(1-\alpha)\right\}.$
This procedure, which we call DA-MCS-marg, yields an MCS with marginal coverage $1-\alpha$ under mild conditions, as formally stated in the following.
Proposition A.1.
Let $M\in\mathds{N}$ be arbitrary and suppose
$\limsup_{n\to\infty}\max_{1\leq m\leq M}\mathds{E}[|L(\widehat{f}_{\mathcal{V}_{m},\mathcal{D}_{\text{train}}},\bm{Z})|^{3}\mid\mathcal{D}_{\text{train}}]<\infty.$
Then for any $m^{*}\in\Theta^{*}$,
$\liminf_{n\to\infty}\mathbb{P}(m^{*}\in\widehat{\Theta}_{marg}\mid\mathcal{D}_{\text{train}},\mathcal{V}_{1},\dots,\mathcal{V}_{M})\geq 1-\alpha.$
The proof is given in Appendix B below. The guarantee in Proposition A.1 is marginal: it holds for any fixed $m^{*}\in\Theta^{*}$. It does not, however, control $\mathbb{P}(\Theta^{*}\subseteq\widehat{\Theta}_{\text{marg}})$ unless $\Theta^{*}$ is a singleton. This motivates the uniform MCS next.
A.3 From Marginal to Uniform MCS Construction: the DA-MCS-plug2/DA-MCS-unif procedure
To construct a uniform MCS, a Bonferroni-type correction is necessary to account for the multiple testing problem arising from simultaneously testing all $M$ models. While replacing $\alpha$ by $\alpha/M$ in the definition of $\widehat{\Theta}_{marg}$ would yield a valid uniform MCS, this approach is often overly conservative, especially when $\Theta^{*}$ is a rather small subset of all models. To address this issue, Kim and Ramdas, (2025) propose a pre-screening step to avoid over-conservatism. The method, which they call DA-MCS-plug2 and we adopt as DA-MCS-unif to clearly distinguish it from DA-MCS-marg, proceeds as follows:
  1. 1.
    For each $r\in\{1,\dots,M\}$, run Algorithm A.3 on the first half of the training losses $\{L_{i,m}\colon 1\leq i\leq n_{\text{val}}/2,m\leq M\}$ at level $n_{\text{val}}^{-1/2}$ to obtain statistics $\widetilde{T}_{r}$. Then compute the of models not rejected at this level:
    $\widetilde{M}=\#\left\{r\in\{1,\dots,M\}\colon\;\widetilde{T}_{r}\leq\Phi^{-1}\bigl(1-n_{\text{val}}^{-1/2}\bigr)\right\}.$
  2. 2.
    For each $r\in\{1,\dots,M\}$, run Algorithm A.3 on all validation losses $\{L_{i,m}\colon 1\leq i\leq n_{\text{val}},1\leq m\leq M\}$ to get $\widehat{T}_{1},\dots,\widehat{T}_{M}$, and set
[…]
A.4 Efficient Implementation and Complexity
When implemented exactly as described above, both of the DA-MCS-marg and the DA-MCS-unif procedures have complexity $O(n_{\text{val}}M^{2})$:
  • Each call to Algorithm A.3 requires $O(n_{\text{val}}M)$ time to compute the cumulative losses, $O(M)$ time to find the argmin, and $O(n_{\text{val}})$ time to compute the sample variance and test statistic. This results in a total of $O(n_{\text{val}}M)$ time per call.
  • Because Algorithm A.3 is called respectively $M$ and $2M$ times in total for DA-MCS-marg and DA-MCS-unif ($M$ times for each step), the overall complexity of both procedures is $O(n_{\text{val}}M^{2})$
@@ A.4 Efficient Implementation and Complexity
  • The column sums $\overline{L}_{m}=\sum_{i\in I_{1}}L_{i,m}$ for $1\leq m\leq M$ can be computed once in $O(n_{\text{val}}M)$ time and accessed in $O(1)$ time in each call to Algorithm A.3.
@@
Figure 5 compares the average computation time of the naive $O(n_{\text{val}}M^{2})$ implementation and the fast $O(n_{\text{val}}M)$ implementation of DA-MCS-plug2 on a log-log scale.
@@
A.5 Runtime Comparison of MCS Methods
In Figure 6, we compare the average computation time of different MCS methods. Method names are from the R script provided by the authors of Kim and Ramdas, (2025), and KR-PLG and KR-PLG-UNIF refer to DA-MCS-marg and DA-MCS-unif, respectively. We observe that both methods are significantly faster than competing methods.
@@ Appendix B PROOFS OF THEORETICAL RESULTS
The proofs of our main results rely on those of Kim and Ramdas, (2025) and a technical lemma, required to match their assumptions with our conditional setup and moment condition. To simplify notation, we write $\mathds{E}^{\prime}[\cdot]=\mathds{E}[\cdot\mid\mathcal{D}_{\text{train}},\mathcal{V}_{1},\dots,\mathcal{V}_{M}]$ and $\mathbb{P}^{\prime}(\cdot)=\mathbb{P}(\cdot\mid\mathcal{D}_{\text{train}},\mathcal{V}_{1},\dots,\mathcal{V}_{M})$ throughout this section.
@@
Let $M\in\mathds{N}$ and define $\Delta_{m,k}=L(\widehat{f}_{\mathcal{V}_{m},\mathcal{D}_{\text{train}}},\bm{X})-L(\widehat{f}_{\mathcal{V}_{k},\mathcal{D}_{\text{train}}},\bm{X})$ and $\overline{\Delta}_{m,k}=\Delta_{m,k}-\mathds{E}^{\prime}[\Delta_{m,k}]$. If
@@
$\mathbb{P}^{\prime}(r\in\widehat{\Theta}_{marg}\mid\widehat{m}_{-r})$ $=\mathbb{P}^{\prime}(\widehat{T}_{r}\leq\Phi^{-1}(1-\alpha)\mid\widehat{m}_{-r})$
@@
To show that the supremum on the right is negligible, we apply the Berry-Esseen bound for Student’s statistic of Bentkus and Götze, (1996), conditionally on $\mathcal{D}_{\text{train}},\mathcal{V}_{1},\dots,\mathcal{V}_{M}$, and $\widehat{m}_{-r}$ (which are all independent of $\{L_{i,m}\colon n_{\text{val}}/2<i\leq n_{\text{val}}\}$). This yields
@@
for a universal constant $C\in(0,\infty)$. Lemma B.1 implies that the right-hand side is bounded by $CK/\sqrt{\lfloor\eta n\rfloor}$ for all $n$ large enough. Altogether, this gives
$\mathbb{P}^{\prime}(r\in\widehat{\Theta}_{marg})=\mathds{E}^{\prime}[\mathbb{P}^{\prime}(r\in\widehat{\Theta}_{marg}\mid\widehat{m}_{-r})]\geq 1-\alpha-\frac{CK}{\sqrt{\lfloor\eta n\rfloor}}=1-\alpha-o(1),$
@@
Proof of Proposition 3.2.
Since $M$ is fixed, Lemma B.1 implies that $\mathbb{P}^{\prime}\in\mathcal{P}^{\leq 3}$ for all $n\geq n_{0}$, where the set $\mathcal{P}^{\leq 3}$ is defined in Kim and Ramdas, (2025, Section 5) as all probability measures $P$ satisfying
$\max_{1\leq m,k\leq M}\frac{\mathds{E}_{P}[|\overline{\Delta}_{m,k}|^{3}]}{\mathds{E}_{P}[|\overline{\Delta}_{m,k}|^{2}]^{3/2}}\leq K,$
for some universal constant $K$. Now Proposition 3.2 from our main paper follows from Theorem 5.1 of (Kim and Ramdas,, 2025). ∎
@@ Appendix C IMPLEMENTATION DETAILS
To ensure both reproducibility and impact, we bundle the proposed methods into a Python package named vinesforests, whose estimators follow the scikit-learn (Pedregosa et al.,, 2011) API (fit, predict, score, score_samples, etc.) and are compatible with numpy (Harris et al.,, 2020) arrays and, where indicated, pandas (The pandas development team,, 2025) DataFrames. As part of the supplementary material, we provide a zip file containing the following file structure:
@@
    • The package exposes two estimator families and their ensemble counterparts: VineDensity, VineRegressor, VineForestDensity, and VineForestRegressor.
    • It provides wrappers KrausVineDensity and KrausVineRegressor around R implementations to reproduce baselines, as well as the following utilities:
        mcs.py — Implementation of the MCS procedures described in Appendix B.
        utils.py — Auxiliary routines for factor expansion and scoring.
        data.py — Dataset loaders used in the experiments.
  • tests/ — Basic unit tests for core functionalities.
  • notebooks/ — Jupyter notebooks allowing to fully reproduce our experiments and analysis.
    • density.ipynb, regression.ipynb, dissmann_in_mcs.ipynb, and analyze_results.ipynb — For running and analyzing the main experiments.
    • mcs.R and argmin_inference.ipynb — For comparing MCS methods and benchmarking the Python implementation.
@@ C.1 Single vine estimators
The class VineBase, which derives from sklearn.base.BaseEstimator, centralizes preprocessing and vine–copula fitting using pyvinecopulib (Vatter and Nagler,, 2025). Inputs may be numpy arrays or pandas DataFrames containing numeric columns, ordered categoricals, and unordered categoricals. Unordered categoricals are expanded to dummy variables via a map $x\mapsto\{\mathbf{1}(x=\ell_{j})\}_{j>1}$, with dummies represented as ordered categoricals taking values $\{0,1\}$. Note that, for both VineBase and its subclasses VineDensity and VineRegressor, expensive computations are batch vectorized and processed in chunks of size batch_size to trade memory for throughput.
@@
  • If a structure is not specified, pyvinecopulib performs structure selection using Dissmann et al., (2013) by default. Otherwise, it needs to be an instance of pyvinecopulib.RVineStructure (e.g., generated via pyvinecopulib.RVineStructure.simulate(d)).
@@ C.2 Forest estimators (ensemble level)
Similarly as in the single vine case, a class VineForestBase deriving from sklearn.base.BaseEstimator implements structure randomization, bootstrap resampling, validation splitting, survivor (either via the MCS or simply those better than Dissmann et al., (2013) in the validation set) selection, and parallel fitting and prediction averaging. It uses a base learner, which can be either VineDensity or VineRegressor, and has two subclasses VineForestDensity and VineForestRegressor. After fitting the default estimator using (Dissmann et al.,, 2013) for structure selection, for a requested $M$ base learners, VineForestBase proceeds as follows:
@@
    Selects survivors either by keeping all candidates strictly improving over the default or by running an MCS procedure at level $\alpha$ (da_mcs_marg and da_mcs_unif for marginal and uniform coverage respectively). Optionally best_only retains the single best survivor.
@@
    Refits all survivors on the full training set and averages predictions at inference time (mean of log-densities for VineForestDensity; averaging conditional weights for VineForestRegressor).
@@
  • Parallelism uses joblib (Joblib Development Team,, 2025). When n_jobs exceeds the number of survivors, the code increases the pyvinecopulib num_threads of each base learner to utilize all cores.
  • The selection routines da_mcs_marg and da_mcs_unif (module mcs.py) implement the procedures described in Section A.2 and Section A.3, with the refinement described in Section A.4 to make it scale linearly with the number of candidates.
@@ C.3 Other Components
Utilities and Datasets
The file utils.py offers: (i) expand_factors for categorical handling; (ii) crps_from_quantiles implementing Simpson–rule CRPS from predictive quantiles. As for the file data.py, it contains lightweight loaders for the UCI/OpenML datasets used in the paper and applies standardization where appropriate; when categorical variables are present (e.g., Bike Sharing), ordered categoricals are preserved for compatibility with VineRegressor.
@@ D.1 Results for Median Regression
In Table 4, we report the average mean absolute error (MAE) on test data for median regressions using the same experimental setup as in our main paper. We observe similar trends as for mean regression and probabilistic forecasting. Our random search methods outperform the greedy Dissmann and Kraus algorithms, with RS-E generally performing best.
@@ D.2 Runtime Comparison on Other Datasets
In Figure 7, we show the relative CPU time of our random search methods compared to Dissmann on all data sets. As expected, the CPU time grows roughly linearly with $M$ in training, irrespective of the random search method or the task. On the other hand, there is no increase in inference time for RS-B, since only one model is used for prediction. For RS-E, the inference time can grow with $M$, since some fraction of the candidates are retained in the MCS and used for prediction. However, this increase is not systematic, as there can sometimes be a single (or a few) model(s) in the MCS, irrespective of $M$.
In Table 5, we report the median CPU time for training and inference using a single vine on all data sets. Except for the California data set, training and inference using a single vine is very fast, taking at most around one second for training and less than 0.5 seconds for inference on a single CPU core. Note that, for inference, the reported CPU time is for predicting the mean, median, and 101 quantiles (from 0% to 100% with a step of 1%) of the conditional distribution of the response given the features at all test points. With around 20’000 data points, the California data set is the largest one considered in our experiments, and for such sizes we would need to optimize how pair-copulas are fitted to further reduce the computational cost. This can be done, for instance, by using a smaller TLL grid size initially (i.e., before applying the MCS), and then re-fitting the selected candidates with a larger grid size.

### v2

@@ 5 Conclusion
We propose a simple and yet effective method for learning vine copula structures based on random search and model confidence sets. Our experiments on real-world data sets demonstrate that the new method outperforms existing greedy heuristics in both density estimation and regression tasks, often by a large margin.
A limitation of the method is the additional computational cost compared to greedy heuristics. However, this cost is modest if the dimension is not too high and easily parallelizable, making it a worthwhile trade-off for the significant performance gains achieved. In truly high-dimensional problems, however, any structure learning method must be combined with additional sparsity-inducing mechanisms, such as truncation, regularization, or variable selection. While beyond the scope of the current work, designing more sophisticated sampling methods that incorporate such mechanisms is an important and promising direction for future research.
Acknowledgements.
We thank Johanna F. Ziegel for pointing us to the literature on model confidence sets. This work was partially supported by the Hasler Stiftung under grant number 2025-05-01-519.
@@ Appendix A DETAILS ON THE MCS ALGORITHM
In the following, we recall the relevant details of the MCS algorithms given by kim2025locally to keep our paper self-contained. Recall the notation of our main paper, specifically Algorithm˜2: There are $M$ candidate models $\{\mathcal{V}_{m}\}_{m=1}^{M}$ and we have already computed validation losses $L_{i,m}$ for $1\leq i\leq n_{\text{val}}$ and $1\leq m\leq M$. For ease of notation, define the loss differences $\Delta_{m,k}^{(i)}=L_{i,m}-L_{i,k}$ for $1\leq i\leq n_{\text{val}}$ and $1\leq m,k\leq M$.
A.1 The DA Test Statistic and DA-MCS-marg Construction
We first describe the discrete argmin (DA) hypothesis test introduced by kim2025locally which serves as the foundation for constructing our MCS. The basic algorithm to compute the test statistic from a set of loss observations is given in Algorithm˜A.3.
@@ A.1 The DA Test Statistic and DA-MCS-marg Construction
To obtain an MCS with marginal coverage guarantees, one can simply invert the DA test: compute $\widehat{T}_{r}$ for all $r=1,\dots,M$ by running Algorithm˜A.3 and define
$\widehat{\Theta}=\left\{r\in\{1,\dots,M\}\colon\;\widehat{T}_{r}\leq\Phi^{-1}(1-\alpha)\right\}.$
This procedure, which we call DA-MCS-marg, yields an MCS with marginal coverage $1-\alpha$ under mild conditions, as formally stated in Proposition˜3.2.
A.2 Efficient Implementation and Complexity
When implemented exactly as described above, the DA-MCS-marg procedure has complexity $O(n_{\text{val}}M^{2})$:
  • Each call to Algorithm˜A.3 requires $O(n_{\text{val}}M)$ time to compute the cumulative losses, $O(M)$ time to find the argmin, and $O(n_{\text{val}})$ time to compute the sample variance and test statistic. This results in a total of $O(n_{\text{val}}M)$ time per call.
  • Because Algorithm˜A.3 is called $M$ times in total for DA-MCS-marg, the overall complexity is $O(n_{\text{val}}M^{2})$
@@ A.2 Efficient Implementation and Complexity
  • The column sums $\overline{L}_{m}=\sum_{i\in I_{1}}L_{i,m}$ for $1\leq m\leq M$ can be computed once in $O(n_{\text{val}}M)$ time and accessed in $O(1)$ time in each call to Algorithm˜A.3.
@@
Figure˜5 compares the average computation time of the naive $O(n_{\text{val}}M^{2})$ implementation and the fast $O(n_{\text{val}}M)$ implementation of DA-MCS-marg (as well as a variant for obtaining uniform MCS) on a log-log scale.
@@
A.3 Runtime Comparison of MCS Methods
In Figure˜6, we compare the average computation time of different MCS methods. Method names are from the R script provided by the authors of kim2025locally; KR-PLG refers to DA-MCS-marg. We observe that the method is significantly faster than competing methods.
@@ Appendix B PROOFS OF THEORETICAL RESULTS
The proof uses similar arguments as kim2025locally, but requires some technical modifications to adapt to our setting. We start with a small lemma simplifying the moment condition required by kim2025locally. To simplify notation, we write $\mathds{E}^{\prime}[\cdot]=\mathds{E}[\cdot\mid\mathcal{D}_{\text{train}}]$ and $\mathbb{P}^{\prime}(\cdot)=\mathbb{P}(\cdot\mid\mathcal{D}_{\text{train}})$ throughout this section.
@@
Let $M\in\mathds{N}$ and define $\Delta_{m,k}=L(\widehat{f}_{\mathcal{V}_{m},\mathcal{D}_{\text{train}}},\bm{Z})-L(\widehat{f}_{\mathcal{V}_{k},\mathcal{D}_{\text{train}}},\bm{Z})$ and $\overline{\Delta}_{m,k}=\Delta_{m,k}-\mathds{E}^{\prime}[\Delta_{m,k}]$. If
@@
$\mathbb{P}^{\prime}(r\in\widehat{\Theta}\mid\widehat{m}_{-r})$ $=\mathbb{P}^{\prime}(\widehat{T}_{r}\leq\Phi^{-1}(1-\alpha)\mid\widehat{m}_{-r})$
@@
To show that the supremum on the right is negligible, we apply the Berry-Esseen bound for Student’s statistic of bentkus1996berry, conditionally on $\mathcal{D}_{\text{train}},\mathcal{V}_{1},\dots,\mathcal{V}_{M}$, and $\widehat{m}_{-r}$ (which are all independent of $\{L_{i,m}\colon n_{\text{val}}/2<i\leq n_{\text{val}}\}$). This yields
@@
for a universal constant $C\in(0,\infty)$. Lemma˜B.1 implies that the right-hand side is bounded by $CK/\sqrt{\lfloor\eta n\rfloor}$ for all $n$ large enough. Altogether, this gives
$\mathbb{P}^{\prime}(r\in\widehat{\Theta})=\mathds{E}^{\prime}[\mathbb{P}^{\prime}(r\in\widehat{\Theta}\mid\widehat{m}_{-r})]\geq 1-\alpha-\frac{CK}{\sqrt{\lfloor\eta n\rfloor}}=1-\alpha-o(1),$
@@ Appendix C IMPLEMENTATION DETAILS
To ensure both reproducibility and impact, we bundle the proposed methods into a Python package named vinesforests, whose estimators follow the scikit-learn (scikit-learn) API (fit, predict, score, score_samples, etc.) and are compatible with numpy (numpy) arrays and, where indicated, pandas (pandas) DataFrames. As part of the supplementary material, we provide a zip file containing the following file structure:
@@
    • vine: This module exposes two estimator classes, VineDensity and VineRegressor, which implement single vine density estimation and regression, respectively.
    • forest: This module exposes two estimator classes, VineForestDensity and VineForestRegressor, which implement the ensemble methods for density estimation and regression, respectively.
    • benchmark: This module exposes two estimator classes, KrausVineDensity and KrausVineRegressor, which are wrappers around R implementations of the baselines.
    • experiments: This module contains utilities for running experiments, including a CLI interface.
    • mcs.py — Implementation of the MCS procedures.
  • tests/ — Unit tests.
  • notebooks/ — Scripts and Jupyter notebooks allowing to fully reproduce our experiments and analysis.
@@ C.1 Single vine estimators
The class VineBase, which derives from sklearn.base.BaseEstimator, centralizes preprocessing and vine–copula fitting using pyvinecopulib (pyvinecopulib). Inputs may be numpy arrays or pandas DataFrames containing numeric columns, ordered categoricals, and unordered categoricals. Unordered categoricals are expanded to dummy variables via a map $x\mapsto\{\mathbf{1}(x=\ell_{j})\}_{j>1}$, with dummies represented as ordered categoricals taking values $\{0,1\}$. Note that, for both VineBase and its subclasses VineDensity and VineRegressor, expensive computations are batch vectorized and processed in chunks of size batch_size to trade memory for throughput.
@@
  • If a structure is not specified, pyvinecopulib performs structure selection using Dissmann2013 by default. Otherwise, it needs to be an instance of pyvinecopulib.RVineStructure (e.g., generated via pyvinecopulib.RVineStructure.simulate(d)).
@@ C.2 Forest estimators (ensemble level)
Similarly as in the single vine case, a class VineForestBase deriving from sklearn.base.BaseEstimator implements structure randomization, bootstrap resampling, validation splitting, survivor (either via the MCS or simply those better than Dissmann2013 in the validation set) selection, and parallel fitting and prediction averaging. It uses a base learner, which can be either VineDensity or VineRegressor, and has two subclasses VineForestDensity and VineForestRegressor. After fitting the default estimator using (Dissmann2013) for structure selection, for a requested $M$ base learners, VineForestBase proceeds as follows:
@@
    Selects survivors either by keeping all candidates strictly improving over the default or by running the MCS procedure at level $\alpha$ (da_mcs_marg). Optionally best_only retains the single best survivor.
@@
    Refits all survivors on the full training set and averages predictions at inference time (log of mean densities for VineForestDensity; averaging conditional weights for VineForestRegressor).
@@
  • Parallelism uses joblib (joblib). When n_jobs exceeds the number of survivors, the code increases the pyvinecopulib num_threads of each base learner to utilize all cores.
  • The selection routine da_mcs_marg (module mcs.py) implement the procedures described in Appendix˜A, with the refinement described in Section˜A.2 to make it scale linearly with the number of candidates.
@@ C.3 Other Components
Experiments
The file utils.py offers: (i) expand_factors for categorical handling; (ii) crps_from_quantiles implementing Simpson–rule CRPS from predictive quantiles. As for the file real_data.py, it contains lightweight loaders for the UCI/OpenML datasets used in the paper and applies standardization where appropriate; when categorical variables are present (e.g., Bike Sharing), ordered categoricals are preserved for compatibility with VineRegressor. The rest of the module contains a CLI interface to run experiments, which is used in the provided Jupyter notebooks to reproduce results.
@@ D.1 Results for Median Regression
In Table˜4, we report the average mean absolute error (MAE) on test data for median regressions using the same experimental setup as in our main paper. We observe similar trends as for mean regression and probabilistic forecasting. Our random search methods outperform the greedy Dissmann and Kraus algorithms, with RS-E generally performing best.
@@ D.2 Runtime Comparison on Other Datasets
In Figure˜7, we show the relative CPU time of our random search methods compared to Dissmann on all data sets. As expected, the CPU time grows roughly linearly with $M$ in training, irrespective of the random search method or the task. On the other hand, there is no increase in inference time for RS-B, since only one model is used for prediction. For RS-E, the inference time can grow with $M$, since some fraction of the candidates are retained in the MCS and used for prediction. However, this increase is not systematic, as there can sometimes be a single (or a few) model(s) in the MCS, irrespective of $M$.
In Table˜5, we report the median CPU time for training and inference using a single vine on all data sets. Except for the California data set, training and inference using a single vine is very fast, taking at most around one second for training and less than 0.5 seconds for inference on a single CPU core. Note that, for inference, the reported CPU time is for predicting the mean, median, and 101 quantiles (from 0% to 100% with a step of 1%) of the conditional distribution of the response given the features at all test points. With around 20’000 data points, the California data set is the largest one considered in our experiments, and for such sizes we would need to optimize how pair-copulas are fitted to further reduce the computational cost. This can be done, for instance, by using a smaller TLL grid size initially (i.e., before applying the MCS), and then re-fitting the selected candidates with a larger grid size.
@@
D.3 Generating vines by local perturbations
In this appendix, we report additional experiments comparing uniform random search to a locally perturbed version of Dissmann’s algorithm. Specifically, instead of selecting the maximum spanning tree (MST) obtained from the empirical Kendall’s $\tau$ matrix at each tree level, we generate random trees according to distribution
$P(T)\propto\prod_{(j,k)\in T}|\tau_{j,k}|,$
with the MST being the mode. Uniform sampling from this distribution can be done efficiently via Wilson’s loop-erased random walk algorithm (wilson1996generating), implementented e.g. in boost (boost) and made available in the vine selection context by pyvinecopulib (pyvinecopulib).
Table 6: Additional results for the experiment in Table˜1 for a variant of Algorithm˜2 that generates vines by sampling local perturbations of the Dissmann structure.
Energy Concrete Airfoil Wine Ccpp California
RS-B (50) 1.54 (0.28) 7.1 (0.17) 3.11 (0.06) 8.47 (0.27) 6.77 (0.01) 3.56 (0.28)
RS-B (100) 1.17 (0.4) 7.11 (0.16) 3.1 (0.07) 8.47 (0.27) 6.77 (0.01) 3.57 (0.28)
RS-B (500) 0.48 (0.67) 6.99 (0.18) 3.11 (0.06) 8.47 (0.27) 6.76 (0.01) 3.58 (0.28)
RS-E (50) 0.53 (0.37) 6.63 (0.13) 3.06 (0.07) 8.19 (0.33) 6.75 (0.01) 3.51 (0.26)
RS-E (100) 0.19 (0.33) 6.61 (0.12) 3.06 (0.07) 8.09 (0.33) 6.75 (0.01) 3.49 (0.27)
RS-E (500) -0.11 (0.36) 6.57 (0.15) 3.07 (0.07) 7.8 (0.21) 6.74 (0.01) 3.46 (0.27)
Table˜6 reports results for the experiment in Table˜1 of the main text. We see that uniform random search slightly outperforms the local perturbation. From a theoretical standpoint, there is no reason to believe that a greedy approach will find a good structure. In fact, it is easy to simulate data where greedy methods (and perturbed variants) fail catastrophically: construct a vine where the first tree consists of weak dependencies, while the later trees contain strong dependencies and asymmetric relationships. The greedy method will pick a structure that severely violates the simplified vine assumption (see Section 2.3), and many randomly sampled structures lead to better performance. However, alternative local methods may be interesting to explore in future work.